You will be provided with a prepared Blue #1 solution with a concentration of 1.0 x 10 -6 M. Household bleach will be provided with a concentration of 6.0% NaOCl by weight. Determine the molarity of the NaOCl.
Answer #1

6 % NaOCl mean 6 g of NaOCl present in 100 mL of solution

molarity = moles / volume of solution in liter

moles = 6 / molar mass of NaOCl

moles = 6 / 74.44 = 0.081 moles

molarity of NaOCl = 0.081 / 0.1 = 0.81 M